Izzy Ice Cream has the following price and cost information:

   Price per 2-scoop sundae $5.00

Variable cost per sundae:  

Ingredients 1.35

Direct labor 0.45

Overhead 0.20

Fixed cost per month $7,500

Required:

1. Determine Izzy's break-even point in units and sales dollars.

2. Determine how many sundaes must be sold to generate a profit of $15,000.

3. Calculate Izzy's new break-even point for each of the following independent scenarios:

a. Sales price decreases by $0.50.

b. Fixed costs decrease by $300 per month.

c. Variable costs increase by $0.50 per sundae.

4. Based on the original information, how many sundaes must Izzy sell to generate a profit of $40,000, if sales price increases by $0.50 and variable costs increase by $0.30?
Answer #1

The answer has been presented in the supporting sheet. For detailed answer refer to the supporting sheet.

Answer Part 1) Break even point in units = fixed cost / contribution margin per unit = 7500/(5-1.35-45-2) = 2500 Units Break

Part 3) New Break even point if sale price decrease by $.5 = Fixed cost / revised contribution margin = 7500 /( 4.5-1.35-45-2